OpenStudy (katie_rogers101):

HELP ME UNDERSTAND If you have a 30-60-90 triangle: 11. And the length of the shortest leg is 4, what's the length of the hypotenuse?

OpenStudy (anonymous):

ratios of a 30:60:90 right triangle are \[1:\sqrt3:2\] for short leg: long leg: hypotenuse

OpenStudy (katie_rogers101):

so the hypotenuse is 2?

OpenStudy (anonymous):

i can show you why if you like

OpenStudy (anonymous):

oh no

OpenStudy (katie_rogers101):

yes please I have alot of questions like this

OpenStudy (katie_rogers101):

and I need to finish these in like 30 minutes

OpenStudy (anonymous):

the hypotenuse is twice the length of the short leg

OpenStudy (katie_rogers101):

oh

OpenStudy (anonymous):

so if your short leg is 4, then the hypotenuse is ___

OpenStudy (katie_rogers101):

8

OpenStudy (anonymous):

bingo next

OpenStudy (katie_rogers101):

so next is12. Working from #11, what's the length of the other leg?

OpenStudy (anonymous):

the short leg times \(\sqrt3\)

OpenStudy (katie_rogers101):

the short leg is times the sqrt of three?

OpenStudy (anonymous):

i wouldn't use a calculator, they probably just want you to write \(4\sqrt3\)

OpenStudy (anonymous):

lol no dear, the length of the short leg times the square root of three !

OpenStudy (katie_rogers101):

Oh okay, is it okay to write it like 4 sqrt 3

OpenStudy (anonymous):

yes

OpenStudy (katie_rogers101):

13. And the length of the longest leg is 5[SQRT(3)], what's the length of the other leg?

OpenStudy (anonymous):

ratios are \[1:\sqrt3:2\] so you have \[4:4\sqrt3:8\]

OpenStudy (anonymous):

in this case you have the long leg is \(5\sqrt3\) so the short leg is that number divided by \(\sqrt3\)

OpenStudy (anonymous):

in other words, cancel the \(\sqrt3\)

OpenStudy (katie_rogers101):

wait so Im confused

OpenStudy (katie_rogers101):

just 5?

OpenStudy (anonymous):

yes

OpenStudy (katie_rogers101):

oh okay

OpenStudy (anonymous):

once again ratios are \[1:\sqrt3:2\] you have \[5:5\sqrt3:10\]

OpenStudy (anonymous):



OpenStudy (katie_rogers101):

14. Working from #13, what's the length of the hypotenuse? The hypotenuse = 10

OpenStudy (katie_rogers101):

is that right

OpenStudy (katie_rogers101):

15. And the length of the longest leg is 9, what is the length of shortest leg?

OpenStudy (anonymous):

yes to 14

OpenStudy (anonymous):

for 15 the short leg is half of the hypotenuse

OpenStudy (anonymous):

oh oops scratch that

OpenStudy (anonymous):

the long let is 9, the short let is 9 divided by the square root of three lets be careful how to compute that

OpenStudy (katie_rogers101):

would we just do 9/ sqrt 3 ?

OpenStudy (anonymous):

\[\frac{9}{\sqrt3}\] is one answer, but they probably want you to do this \[\frac{9}{\sqrt3}\times \frac{\sqrt3}{\sqrt3}=\frac{9\sqrt3}{3}=3\sqrt3\]

OpenStudy (katie_rogers101):

oh okay

OpenStudy (anonymous):

the last one \(3\sqrt3\) is the same number, just in "simplest radical form"

OpenStudy (katie_rogers101):

16. Working from #15, what is the length of the hypotenuse?

OpenStudy (anonymous):

double the short leg

OpenStudy (katie_rogers101):

so it would be double the shortest leg

OpenStudy (anonymous):

you are getting it i see

OpenStudy (katie_rogers101):

but the shortest leg is 3 sqrt 3

OpenStudy (anonymous):

right multiply that by 2

OpenStudy (katie_rogers101):

6 sqrt 6

OpenStudy (anonymous):

oh no just \(6\sqrt3\) not \(6\sqrt6\)

OpenStudy (katie_rogers101):

oh okay

OpenStudy (anonymous):

on account of \[2\times 3\sqrt3=6\sqrt3\] just like \[2\times 3x=6x\]

OpenStudy (katie_rogers101):

17. With a hypotenuse of 2[SQRT(3)] what is the length of the longest leg?

OpenStudy (katie_rogers101):

ohhh and i see

OpenStudy (anonymous):

the short leg is ...

OpenStudy (katie_rogers101):

the answer is 2

OpenStudy (anonymous):

nope

OpenStudy (katie_rogers101):

dang

OpenStudy (anonymous):

divide the hypotenuse by 2 to get the short leg the long leg is 2

OpenStudy (katie_rogers101):

wait wait

OpenStudy (anonymous):

ok

OpenStudy (katie_rogers101):

so 2/2 is 1

OpenStudy (katie_rogers101):

right is that what im doing?

OpenStudy (katie_rogers101):

this one confuses me

OpenStudy (anonymous):

right lets back up a second cause i made a mistake

OpenStudy (anonymous):

the hypotenuse is \(2\sqrt3\) right?

OpenStudy (katie_rogers101):

yes

OpenStudy (anonymous):

the short leg is therefore half of that

OpenStudy (katie_rogers101):

so the sqrt 3 cancels out right?

OpenStudy (anonymous):

no not dividing by \(\sqrt3\) just divide by 2

OpenStudy (katie_rogers101):

oh so 1

OpenStudy (anonymous):

\[\frac{2\sqrt3}{2}=?\]

OpenStudy (anonymous):

think of it like \[\frac{2x}{2}\]

OpenStudy (katie_rogers101):

im still really confused

OpenStudy (anonymous):

ok i don't want to confuse you lets see if i can explain the short leg is half the hypotenuse whatever the hypotenuse is, divide it by 2 half of \(2\sqrt3\) is \(\sqrt3\)

OpenStudy (anonymous):

just like half of \(2x\) is \(x\)

OpenStudy (katie_rogers101):

Ohhhhokay

OpenStudy (katie_rogers101):

so the short leg is just sqrt 3

OpenStudy (anonymous):

yes, exactly

OpenStudy (anonymous):

now i bet you need the long leg right?

OpenStudy (katie_rogers101):

right

OpenStudy (anonymous):

it is the short leg times \(\sqrt3\) since the short leg is \(\sqrt3\) itself, you need to multiply \[\sqrt3\times \sqrt3\]

OpenStudy (katie_rogers101):

sqrt 9

OpenStudy (katie_rogers101):

I think

OpenStudy (anonymous):

lol yeah but c'mon what is \(\sqrt9\)?

OpenStudy (katie_rogers101):

3